### v2

# Reconciling Reality through Simulation: A Real-to-Sim-to-Real Approach for Robust Manipulation

Abstract
Imitation learning methods need significant human supervision to learn policies robust to changes in object poses, physical disturbances, and visual distractors. Reinforcement learning, on the other hand, can explore the environment autonomously to learn robust behaviors but may require impractical amounts of unsafe real-world data collection. To learn performant, robust policies without the burden of unsafe real-world data collection or extensive human supervision, we propose RialTo, a system for robustifying real-world imitation learning policies via reinforcement learning in “digital twin” simulation environments constructed on the fly from small amounts of real-world data. To enable this real-to-sim-to-real pipeline, RialTo proposes an easy-to-use interface for quickly scanning and constructing digital twins of real-world environments. We also introduce a novel “inverse distillation” procedure for bringing real-world demonstrations into simulated environments for efficient fine-tuning, with minimal human intervention and engineering required. We evaluate RialTo across a variety of robotic manipulation problems in the real world, such as robustly stacking dishes on a rack, placing books on a shelf, and six other tasks. RialTo increases (over 67%) in policy robustness without requiring extensive human data collection. Project website and code at https://real-to-sim-to-real.github.io/RialTo/.
footnotetext: * Equal advising
texs/intro2
I Related Work
Learning Visuomotor Control from Demonstrations: Behavior cloning (BC) of expert trajectories can effectively acquire robot control policies that operate in the real world [19, 13, 68, 6, 20, 37]. While several works have used BC to learn performant policies from small to moderately-sized datasets [13, 68, 37], performance tends to drop when the policy must generalize to variations in scene layouts and appearance. Techniques for improving BC often require much larger-scale data collection [6, 54], raising scalability concerns. Other techniques support generalization with intermediate representations  [20] and leverage generative models to add visual distractors [67, 36]. These can improve robustness to visual distractors but do not address physical or dynamic disturbances, as these require producing actions not present in the data.
Fine-tuning Imitation with RL and Improving RL with Demonstrations: Reinforcement learning has been used to improve the performance of models originally trained with imitation learning. RL has exploded in its capacity for fine-tuning LLMs [47] and image generation models [4], learning rewards from human feedback [14]. In robotics, prior work has explored techniques such as offline RL [66, 45, 31], learning world models [40, 18], and online fine-tuning in the real world [2, 3, 22, 66]. Expert demonstrations have also been used to bootstrap exploration and policy learning with RL [27, 26, 52, 70]. We similarly combine imitation and RL to guide exploration in sparse reward settings. However, our pipeline showcases how demonstrations additionally benefit RL by biasing policies toward physically plausible solutions that compensate for imperfect physics simulation.
Sim-to-real policy transfer: RL in simulation has been used to synthesize impressive control policies in a variety of domains such as locomotion [38, 32, 30], dexterous in-hand manipulation [11, 12, 1, 23], and drone flight [57]. Many simulation-based RL methods leverage some form of domain randomization [61, 49], system identification [25, 59], or improved simulator visuals [53, 24] to reduce the simulation-to-reality (sim-to-real) domain gap. Prior work has also shown the benefit of “teacher-student” distillation [12, 30, 56, 9], wherein privileged “teacher” policies learned quickly with RL are distilled into “student” policies that operate on sensor observations. To acquire transferable controllers, we similarly leverage GPU-accelerated simulation, teacher-student training, and domain randomization across parallel environments. However, we address the more challenging scenario of household manipulation, which is characterized by richer visual scenes, and minimize the necessary engineering effort by relying on sparse rewards. We also simplify sim-to-real by training on digital twin assets and co-training with real data [63].
Real-to-sim transfer of scenes: Designing realistic simulation environments has been studied from the perspective of synthesizing digital assets that reflect real objects. Prior work has used tools from 3D reconstruction [28] and inverse graphics [10] for creating digital twins, and such real-to-sim pipelines have been used for both rigid and deformable [58] objects. These approaches are all compatible with our system and could be used to automate real-to-sim scene transfer and reduce human effort. Our work similarly leverages advancements in 3D vision [60] for reconstructing object geometry, but we also introduce an easy-to-use GUI for building a URDF/USD with accurate articulations. Furthermore, our GUI could be used to improve the aforementioned methods by making it easier to collect a large dataset of human-annotated articulated scenes. The accuracy of the simulator could be improved further combining our GUI with the latest system identification research [39, 33].
Real-to-sim-to-real transfer: Prior work has used NeRF [41] and other 3D reconstruction techniques to create realistic scene representations for improving manipulation  [69], navigation [16, 8] and locomotion [7]. These works, however, only use the visual component of the synthetic scene and do not involve any physical interaction with a reconstructed geometry. As a result, these systems cannot adjust to environmental changes beyond visual distractions. For instance, different grasp poses may require different placements, and a policy cannot discover these novel behaviors without physically interacting with the environment during training. A limited number of works have learned policies that interact with the reconstructed environments, but they either simplify the reconstructed shapes [35] or are limited to simple grasp motions [64].
II RialTo: A Real-to-Sim-to-Real System for Robust Robotic Manipulation
Figure 1: Overview of the real-to-sim pipeline for transfering scenes to the simulator. The first stage consists of scanning the environment, using off-the-shelf tools such as NeRFStudio, ARCode, or Polycam. Each has its strengths and weaknesses and should be used appropriately (see Appendix XI for recommendations). The second stage consists of uploading the reconstructed scene into RialTo’s GUI where the user can cut the mesh, specify joints, and organize the scene as desired. Once complete, the scene can be downloaded as a USD asset, which can be directly imported into the simulator.
II-A System Overview
Our goal is to obtain a control policy that maps real-world sensory observations to robot actions. We only assume access to a small set of demonstrations $(\sim 15)$ containing (observation, action) trajectories collected by an expert, although in principle RialTo can also be used to robustify large, expressive pretrained models as well. Our approach robustifies real-world imitation learning policies using simulation-based RL to make learned controllers robust to disturbances and distractors not present in the demos. The proposed pipeline, RialTo, achieves this with four main steps (Fig LABEL:fig:teaser):
  1. 1.
    We construct geometrically, visually, and kinematically accurate simulation environments from real-world image capture. We leverage 3D reconstruction tools and develop an easy-to-use graphical interface for adding articulations and physical properties.
  2. 2.
    We obtain a set of successful trajectories containing privileged information (such as Lagrangian state, e.g. object and joint poses) in simulation. We propose an “inverse distillation” algorithm to transfer a policy learned from real-world demonstrations to create a dataset of trajectories (i.e., demos) in the simulation environment.
  3. 3.
    The synthesized simulation demos bootstrap efficient fine-tuning with RL in simulation using an easy-to-design sparse reward function and low-dimensional state space, with added randomization to make the policy robust to environmental variations.
  4. 4.
    The learned policy is transferred to reality by distilling a state-based simulation policy into a policy operating from raw sensor observations available in the real world [9, 12]. During distillation, we also co-trained with the original real-world demonstrations to capitalize on the combined benefits of simulation-based robustification and in-domain real-world data.
The following sections describe each component in detail, along with a full system overview in Fig LABEL:fig:teaser.
Figure 2: Inverse distillation & RL fine-tuning. We introduce a novel procedure for going from point cloud-based policies trained from real-world demonstrations $\mathcal{D}_{\text{real}}$ to a robust privileged state-based policy in simulation. 1) Train a vision-based policy with supervised learning on $\mathcal{D}_{\text{real}}$ 2) Rollout the vision-based policy on the simulation rendered point clouds and collect a set of 15 privileged demonstrations with object poses, $\mathcal{D}_{\text{sim}}$ 3) Train a robust state-based policy with RL and a sparse reward, adding a BC loss fitting $\mathcal{D}_{\text{sim}}$ to bias exploration and set a prior on real-world-transferable policies.
II-B Real-to-Sim Transfer for Scalable Scene Generation
The first step of RialTo is to construct geometrically, visually, and kinematically realistic simulated scenes for policy training. This requires (i) generating accurate textured 3D geometry from real-world images and (ii) specifying articulations and physical parameters. For geometry reconstruction, we use existing off-the-shelf 3-D reconstruction techniques. Our pipeline is agnostic to the particular method used, and we have verified the approach with a variety of scanning apps (e.g., Polycam [50] and ARCode [15]) and 3D reconstruction pipelines [60, 43], each of which convert a set of multi-view 2D images (or a video) into a textured 3D mesh. The raw mesh denoted $\mathcal{G}$, is typically exported as a single globally-unified geometry, which is unsuitable for direct policy learning. Scene objects are not separated and the kinematics of objects with internal joints are not reflected. Physical parameters like mass and friction are also required and unspecified. We therefore further process the raw mesh $\mathcal{G}$ into a set of separate bodies/links $\{\mathcal{G}_{i}\}_{i=1}^{M}$ with kinematic relations $\mathcal{K}$ and physical parameters $\mathcal{P}$.
While there are various automated techniques for automatically segmenting and adding articulations to meshes [28], in this work, we take a simple human-centric approach. We offer a simple graphical interface for humans to quickly separate meshes and add articulations (see Fig. 1). Our GUI allows users to upload their own meshes and drag/drop, reposition, and reorient them in the global scene. Users can then separate meshes and add joints between different mesh elements, allowing objects like drawers, fridges, and cabinets to be scanned and processed. Importantly, our interface is lightweight, intuitive, and requires minimal domain-specific knowledge. We conducted a study (Section V) evaluating six non-expert users’ experiences with the GUI and found they could scan complex scenes and populate them with a couple of articulated objects in under 15 minutes of active interaction time. Examples of real-world environments with their corresponding digital twins are shown in Fig 3 and Appendix Fig. 15.
The next question is —how do we infer the physics parameters that faithfully replicate the real world? While accurately identifying physical parameters is possible, this can be challenging without considerable interaction [5, 65]. While adapting to dynamics variations is an important direction for future work, in this system we set a single default value for mass and friction uniformly across objects and compensate for the sim-to-real gap to actual real-world values by constraining the learned policy to be close to a small number of real-world demonstrations as discussed in Section II-C.
This procedure produces a scene $\mathcal{S}=\{\{\mathcal{G}_{i}\}_{i=1}^{M},\mathcal{K},\mathcal{P}\}$ represented in a USD/URDF file that references the separated meshes and their respective geometric ($\mathcal{G}_{i}\}_{i=1}^{M}$), kinematics ($\mathcal{K}$) and physical parameters ($\mathcal{P}$). This environment can subsequently be used for large-scale policy robustification in simulation.
II-C Robustifying Real-World Imitation Learning Policies in Simulation
Given the simulation environment generated in Section II-B, the next step in RialTo involves learning a robust policy in simulation that can solve desired tasks from a wide variety of configurations and environmental conditions. While this can be done by training policies from scratch in simulation, this is often a prohibitively slow process, requiring considerable manual engineering. Instead, we will adopt a fine-tuning-based approach, using reinforcement learning in simulation to fine-tune a policy initialized from a small number of expert demonstrations collected in the real world. Since training RL directly from visual observations is challenging, we would ideally like to finetune simulation policies that are based on a privileged Lagrangian state. However, real-world demonstrations do not have access to the low-level state information in the environment. To enable the bootstrapping of RL finetuning in simulation from a privileged state using real-world demonstrations, we introduce a novel “inverse distillation” (Section II-C1) procedure that is able to take real-world demonstrations with only raw sensor observations and actions and transfer them to simulation demonstrations, complete with low-level privileged state information. These privileged information demonstrations can then be used to instantiate an efficient RL-based fine-tuning procedure (Section II-C2) in simulation to massively improve policy robustness.
II-C1 Inverse-distillation from Real-to-Sim for Privileged Policy Transfer
We assume a human provides a small number of demonstrations in the real world $\mathcal{D}_{\text{real}}=\{(o_{1}^{i},a_{1}^{i}),\dots,(o_{H}^{i},a_{H}^{i})\}_{i=1}^{N}$, where trajectories contain observations $o$ (3D point clouds) and actions $a$ (delta end-effector pose). Considering that simulation-based RL fine-tuning is far more efficient and performant when operating from a compact state representation [30, 11] (see Section IV-C) and we wish to use real-world human demonstrations to avoid the difficulties with running RL from scratch (see Section IV-B), we want to transfer our observation-action demonstrations from the real world to simulation in a way that allows for subsequent RL fine-tuning in simulation from compact state-based representations. This presents a challenge because we do not have an explicit state estimation system that provides a Lagrangian state for the collected demonstrations in the real world. We instead introduce a procedure, called “inverse-distillation”, for converting our real-world set of demonstrations into a set of trajectories in simulation that are paired with privileged low-level state information.
Given the demonstrations $\mathcal{D}_{\text{real}}$, we can naturally train a policy $\pi_{\text{real}}(a|o)$ on this dataset via imitation learning. “Inverse distillation” involves executing this perception-based learned policy $\pi_{\text{real}}(a|o)$ in simulation, based on simulated sensor observations $o$, to collect a dataset $\mathcal{D}_{\text{sim}}=\{(o_{1}^{i},a_{1}^{i},s_{1}^{i})\dots,(o_{H}^{i},a_{H}^{i},s_{H}^{i})\}_{i=1}^{M}$ of successful trajectories which contain privileged state information $s_{t}^{i}$. The key insight here is that while we do not have access to the Lagrangian state in the real-world demonstrations when a learned real-world imitation policy is executed from perceptual inputs in simulation, low-level privileged Lagrangian state information can naturally be collected from the simulation as well since the pairing between perceptual observations and Lagrangian state is known apriori in simulation. Since the goal is to improve beyond the real-world imitation policy $\pi_{\text{real}}(a|o)$, we can then perform RL fine-tuning, incorporating the privileged demonstration dataset $\mathcal{D}_{\text{sim}}$ into the training process, as discussed in the following subsection.
II-C2 Reinforcement Learning Fine-tuning in Simulation
Given the privileged information dataset $\mathcal{D}_{\text{sim}}$, and the constructed simulation environment the goal is to learn a robust policy $\pi^{*}_{\text{sim}}(a|s)$ using reinforcement learning. There are two key challenges in doing so in a scalable way: (1) resolving exploration challenges with minimal reward engineering, and (2) ensuring the policy learns behaviors that will transfer to the real world. ‘ We find that both challenges can be addressed by a simple demonstration augmented reinforcement learning procedure [56, 44, 52], using the Lagrangian state-based dataset $\mathcal{D}_{\text{sim}}$. To avoid reward engineering, we define a simple reward function that detects if the scene is in a desired goal state (detailed sparse reward functions used in each task in Appendix VII). We build on the proximal policy optimization [55] algorithm with the addition of an imitation learning loss as follows (where $\hat{A}_{t}$ is the estimator of the advantage function at step $t$ [55], and $V_{\phi}$ is the learned value function):
$\begin{split}\max_{\theta,\phi}\alpha\sum_{(s_{t},a_{t},r_{t})\in\tau_{\pi_{\theta_{\text{old}}}}}\text{min}(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}(a_{t}|s_{t})}}\hat{A}_{t},\\ \text{clip}(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}(a_{t}|s_{t})}},1-\epsilon,1+\epsilon)\hat{A}_{t})\\ +\beta\sum_{(s_{t},V_{t}^{\text{targ}})\in\tau_{\pi_{\theta_{\text{old}}}}}(V_{\phi}(s_{t})-V_{t}^{\text{targ}})^{2}\\ +\gamma\sum_{(s_{i},a_{i})\in\mathcal{D}_{\text{sim}}}\frac{\pi_{\theta}(a_{i}|s_{i})}{\sum_{a_{c}}\pi_{\theta}(a_{c}|s_{i})}\end{split}$ (1)
In addition to mitigating issues associated with exploration [44, 52], leveraging the additional imitation learning term in the objective helps bias the policy toward physically plausible, safe solutions that improve transfer of behaviors to reality. During this process, we can train the policy for robustness by randomizing initial robot/object/goal poses. Appendix VII contains complete details of our training procedure. The result is a robust policy $\pi^{*}_{\text{sim}}(a|s)$ operating from Lagrangian state that is successful from a wide variety of configurations and environmental conditions.
II-D Teacher-Student Distillation with Co-Training on Real-World Data for Sim-to-Real Transfer
In previous sections, we described a method for efficiently learning a robust policy $\pi^{*}_{\text{sim}}(a|s)$ in simulation using privileged state information. However, in the real world, this privileged information is unavailable. Policy deployment requires operating directly from sensory observations (such as point clouds) in the environment. To achieve this, we build on the framework of teacher-student distillation (with interactive DAgger labeling)[54, 12] where the privileged information policy $\pi^{*}_{\text{sim}}(a|s)$ serves as a teacher and the perceptual policy $\pi^{*}_{\text{real}}(a|o)$ is the studentFor the sake of this work, we will assume that the optimal actions for the student and teacher coincide, and there are no information gathering specific challenges induced by partial observability [56]. Since there is inevitable domain shift between simulation and real domains, this training procedure can be further augmented by co-training the distillation objective with a mix of the original real-world demonstration data $\mathcal{D}_{\text{real}}$ and simulation data drawn from $\pi^{*}_{\text{sim}}(a|s)$ (via the DAgger objective [12]). This results in the following co-training objective for teacher-student policy learning:
$\begin{split}\max_{\theta}\alpha\sum_{(s_{i},o_{i},a_{i})\sim\tau_{\pi_{\theta}}}\frac{\pi_{\theta}(\pi_{\text{teacher}}(s_{i})|o_{i})}{\sum_{a_{c}}\pi_{\theta}(a_{c}|o_{i})}\\ +\beta\sum_{(o_{i},a_{i})\in\mathcal{D}_{\text{real}}}\frac{\pi_{\theta}(a_{i}|o_{i})}{\sum_{a_{c}}\pi_{\theta}(a_{c}|o_{i})}\end{split}$ (2)
Here the first term corresponds to DAgger training in simulation, while the second term co-trains on real-world expert data. This allows the policy to take advantage of small amounts of high-quality real-world data to bridge the perceptual gap between simulation and real-world scenes and improve generalization compared to only using the data from simulation. We empirically demonstrate (Section II-D) that this significantly increases the resulting success rate in the real world. On a practical note, we refer the reader to Appendix VIII for additional details on the student-teacher training scheme that enables it to be successful in the proposed problem setting.
Figure 3: We depict the six tasks used to evaluate RialTo. From top to bottom, we first show the original environment where we collect the demonstrations, second the simulated environment, third the environment where we do our final evaluation containing clutter and disturbances, and fourth the task randomization overview each shaded area corresponds to an approximation of how much randomization each object/robot have.
III Experimental Evaluation
Our experiments are designed to answer the following questions about RialTo: (a) Does RialTo provide real-world policies robust to variations in configurations, appearance, and disturbances? (b) Does co-training policies with real-world data benefit real-world evaluation performance? (c) Is the real-to-sim transfer of scenes and policies necessary for training efficiency and the resulting performance? (d) Does RialTo scale up to more in-the-wild scenes?
To answer these questions, we evaluate RialTo in eight different tasks, shown in Figure 3 and 7. These include 6-DoF grasping and reorientation of free objects (book on a shelf, plate on a rack, mug on a shelf) and 6-DoF grasping and interacting with articulated objects (drawer and cabinet) on a tabletop and opening a toaster, plate on a rack, putting a cup in the trash in more uncontrolled scenes. More details on the tasks such as their sparse reward functions and randomization setups are presented in Appendix VII. For each task, we consider three different disturbance levels in increasing order of difficulty (see Appendix VII for more details):
  1. 1.
    Randomizing object poses: at the beginning of each episode we randomize the object and/or robot poses.
  2. 2.
    Adding visual distractors: at the beginning of each episode we also add visual distractors in a cluttered way.
  3. 3.
    Applying physical disturbances: we apply physical disturbances throughout the episode rollout. We change the pose of the object being manipulated or the target location where the object needs to be placed, close the drawer/toaster/cabinet being manipulated, and move the robot base when possible.
We conduct our experiments on a Franka Panda arm with the default parallel jaw gripper, using 6 DoF Cartesian end effector position control. For perceptual inputs, we obtain 3D point cloud observations from a single calibrated depth camera. More details on the hardware setup can be found in Appendix IX. All of the results in the real world are evaluated using the best policy obtained for each method, we report the average across at least 10 rollouts and the bootstrapped standard deviation. Videos of highlights and evaluation runs are available in the website.
Throughout the next sections, we will evaluate RialTo against the following set of baselines and ablations: 1) Imitation learning from 15 and 50 demos (Section III-A); 2) No co-training on real-world data (Section III-B); 3) Co-training on demonstrations in simulation (Section III-B); 4) RialTo from simulation demos (Section III-C2); 5) Learning from an untargeted set of simulated assets (Section III-C1); 6) RialTo without distractors (Section IV-A); 7) RialTo without demos (Section IV-B)
III-A RialTo Learns Robust Policies via Real-to-Sim-to-Real
Figure 4: Comparison of RialTo against imitation learning both from 15 demonstrations. RialTo provides robust policies across tasks and levels of distractions while imitation learning severely suffers when adding distractors and disturbances.
In this section, we aim to understand whether RialTo can solve complex tasks, showing robustness to variations in configurations, disturbances, and distractors. We compare our approach of real-to-sim-to-real RL fine-tuning against a policy trained only on real-world demos via standard imitation learning (BC). We report the results of running RialTo’s pipeline starting from 15 demos collected directly in simulation and co-training with 15 real-world demos during the teacher-student distillation. In Section III-C2 we show a comparison of running RialTo uniquely on real or sim demos.
The results in Figure 4 show RialTo maintains high performance across configuration levels, achieving on average 91% success across tasks for randomizing object poses, 77% with distractors, and 75% with disturbances. On the other hand, the presence of distractors and disturbances severely reduces the performance of pure imitation learning. For instance, when only randomizing the object poses, the BC baseline achieves an average of 25% success rate across tasks. Under more challenging conditions, the BC baseline drops to 11% and 5% overall performance on average for distractors and disturbances, respectively.
Figure LABEL:fig:teaser, 9 and the videos in the website qualitatively show how the resulting policies are robust to many kinds of environment perturbations, including moving the robot, moving the manipulated object and target positions, and adding visual distractors that cause occlusion and distribution shift. The policy rollouts also demonstrate error recovery capabilities, correcting the robot’s behavior in closed loop when, e.g., objects are misaligned or a grasp must be reattempted. This highlights that RialTo provides robustness that does not emerge by purely learning from demonstrations.
Only randomization Distractors Disturbances
BC (15 demos) 10 $\pm$ 9% 0 $\pm$ 0% 0 $\pm$ 0%
BC (50 demos) 40 $\pm$ 15% 30 $\pm$ 16% 20 $\pm$ 13%
RialTo (15 demos) 90 $\pm$ 9% 70 $\pm$ 14% 60 $\pm$ 16%
TABLE I: RialTo and imitation learning on placing a book on the shelf.
We also compare RialTo against behavior cloning with 50 demonstrations to show that the problem is not simply one of slightly more data. Collecting the 50 demonstrations takes a total time of 1 hour and 45 minutes, which is significantly more than the human effort for RialTo for which we collect 15 demos, in 30 minutes, and build the environment in 15 minutes of active time (see Section V). Although more data improves the performance of direct imitation learning from 10% to 40%, 0% to 30%, and 0% to 20% for the three different levels of robustness, the results in Table I show that RialTo achieves approximately 2.5 times higher success rate than pure BC, despite using less than one third the number of demonstrations and taking less than half of the human supervision’s time.
III-B Impact of Co-Training with Real-World Data
Next, we assess the benefits offered by co-training with real-world demonstrations during teacher-student distillation, rather than just purely training policies in simulation. We consider the book on shelf, plate on rack, mug on shelf, and open drawer tasks (the two first being two of the harder tasks with lower overall performance). The results in Figure 5 illustrate that co-training the policy with 15 real-world demonstrations significantly increases real-world performance on some tasks(3.5x and 2x success rate increase for book on shelf and plate on rack with disturbances, when comparing co-training on real-world demos against co-training with sim demos) while keeping the same performance on tasks that already have a small sim-to-real gap. Qualitatively, we observe the co-trained policy is more conservative and safer to execute. For instance, the policy without co-training usually comes very close to the plate or the book, occasionally causing it to fall. The policy with co-training data, however, leaves more space between the hand and the book before grasping, which is closer to the demonstrated behavior. The observation that sim co-training performs significantly worse than real-world co-training, indicates that co-training with real-world demonstrations is helping in reducing the sim-to-real gap for both the visual distribution shift between simulated and real point clouds and the sim-to-real dynamics gap.
Figure 5: Comparison between running RialTo on sim vs real data. The performance on the methods doing co-training with real-world demos is higher than using only simulated demos or no real-world co-training, on the harder tasks( plate on rack and book on shelf), and matches the performance in the easier tasks( open drawer and mug on shelf). Furthermore, starting from real-world or simulated demos does equally well.
III-C Is Real-to-Sim Transfer Necessary?
Figure 6: Comparison between training with RialTo on the reconstruction of the target drawer against training on a set of four drawers from Objaverse[17]. We observe, that RialTo on the real-to-sim asset does significantly better (90% vs 10%) when testing in the real world on the target drawer compared to training on the set of randomized drawers.
III-C1 Real-to-Sim Transfer of Scenes
Instead of reconstructing assets from the target environment, one could train a policy on a diverse set of synthetic assets and hope the model generalizes to the real-world target scene [12, 21, 62]. While this has shown promising results for object-level manipulation, such as in-hand reorientation  [12], it is still an active area of work for scene-level manipulation and rearrangement  [21]. Moreover, such methods require significant effort in creating a dataset of scenes and objects that enables the learned policies to generalize. Acquiring a controller that can act in many scenes is also a more challenging learning problem, requiring longer wall clock time, more compute, and additional engineering effort to train a performant policy on a larger and more diverse training set.
To probe the benefits of RialTo over such a sim-only training pipeline, we compared the performance against a policy trained using only synthetic assets. Using an amount of time effort roughly comparable to what is required from a single user following our real-to-sim approach (see Section V), we collected a set of 4 drawers from the Objaverse dataset (see Figure 6). Although this is small compared to the growing size of 3D object datasets, we found it non-trivial to transfer articulated objects into simulation-ready USDs and we leave it as future work. Given these manually constructed diverse simulation scenes, we then trained a multi-task policy using RialTo from 20 demonstrations to open the 4 drawers. See Appendix VIII-C for the minor modifications to incorporate multi-task policy learning to RialTo.
As shown in Figure 6, when evaluating the real target drawer, the policy trained on multiple drawers only achieves a 10% success rate, much lower than the 90% obtained by the policy trained on the target drawer in simulation. This leads us to conclude that to train a generalist agent, considerably more data and effort are needed as compared to the relatively simple real-to-sim procedure we describe for test time specialization. Moreover, this suggests that for performance on particular deployment environments, targeted generation of simulation environments via real-to-simulation pipelines may be more effective than indiscriminate, diverse procedural scene generation.
III-C2 Real-to-sim transfer of policies
We additionally want to understand the impact of transferring policies from real-world demonstrations in comparison to running the pipeline starting with demos collected directly in simulation. This helps analyze whether instead of collecting demos both in simulation and in the real world (for the co-training) we can simply collect demos in the real world and do all the training with those.
Figure 5 shows the real-world performance of policies trained using RialTo when starting the RL fine-tuning step using real-world demonstrations as explained in II-C1 against using demonstrations provided directly in simulation. We observe that the performance for both cases is very close. These results show that RialTo successfully learns policies with demonstrations from either source of supervision as long as we keep co-training the policies with real-world data in the teacher-student distillation step. Firstly, this indicates that we do not need to collect both demos in sim and real, but we can run RialTo uniquely from the demos in the real world. Furthermore, this flexibility is a strength of our pipeline, as the ease of acquiring different sources of supervision may vary across deployment scenarios – i.e., one could use policies pretrained from large-scale real-world data or obtain data from a simulation-based crowdsourcing platform.
III-D Scaling RialTo to In-the-Wild Environments
Figure 7: We test RialTo on uncontrolled and in-the-wild scenes, and we see we can continue to solve a variety of tasks more robustly than imitation learning techniques.
In this section, we scale up RialTo to more uncontrolled and in-the-wild environments. We test RialTo on three different tasks: open the microwave in a kitchen (also shown in Section III-A), put a cup in the trash, and bring the plate from the sink to the dishrack. We observe that RialTo scales up to these more diverse scenes and continues to perform significantly better than standard imitation learning techniques. In particular, RialTo brings on average a 57% improvement upon standard imitation learning, see Fig 7.
IV Further Analysis and Ablations
Pose Randomization Distractors
RialTo without distractor training 60 $\pm$ 15% 30 $\pm$ 15%
RialTo with distractor training 100 $\pm$ 0% 70 $\pm$ 15%
TABLE II: Real-World performance of policies trained with and without distractors on the task of placing a mug on a shelf.
IV-A Training with Distractors
When performing teacher-student distillation we performed randomization with additional visual distractors to train a more robust policy that succeeds even in visual clutter. We analyze how this affects the final robustness of the learned policy. For the sake of analysis, we consider the performance on the mug on the shelf task. The small size of the mug and its resemblance in shape and size to other daily objects make the visual component of this task particularly challenging when other objects are also present. Our findings in Table II show that adding distractors during training increases the success rate from 30% to 70% when testing the policy in environments with distractors. We also observe a performance improvement in setups with no distractors suggesting that such training also supports better sim-to-real policy transfer.
IV-B Comparison to RL from Scratch
Open Book on Plate on Mug on Open
toaster shelf rack shelf drawer
RL from scratch with 0 demos 62 $\pm$ 2% 0 $\pm$ 0% 2 $\pm$ 0% 0 $\pm$ 0% 0 $\pm$ 0%
RL fine-tuning from 15 real demos 91 $\pm$ 1% 90 $\pm$ 1% 81 $\pm$ 2% 81 $\pm$ 2% 96 $\pm$ 1%
RL fine-tuning from 15 sim demos 96 $\pm$ 1% 89 $\pm$ 1% 82 $\pm$ 2% 82 $\pm$ 2% 95 $\pm$ 1%
TABLE III: Comparison of training RL from scratch against RL from real and sim demos. RL from sim and real demos seem to be equivalent in most cases, but RL from scratch barely solves the task.
[…]
IV-C RL from Vision
RialTo’s “inverse distillation” procedure to a compact state-space adds some methodological overhead to the system when compared to the possibility of doing RL fine-tuning directly on visual observations. However, as reported in Appendix Fig. 13, on the task of drawer opening, RL from compact states achieves a 96% success rate after 12 hours of wall-clock time, while RL from vision only achieves a 1% success rate after 35 hours. Hence, inverse distilling to state space is necessary because training RL from vision with sparse rewards is prohibitively slow, motivating the methodology outlined in Section II-C1.
V User study
We analyzed the usability of RialTo’s pipeline for bringing real-world scenes to simulation. We ran a user study over 6 people, User 6 being an expert who used the GUI before and Users 1-5 never did any work on simulators before. Each participant was tasked with creating an articulated scene using the provided GUI. More precisely, their task was to: 1) scan a big scene, 2) cut one object, 3) scan and upload a smaller object, and 4) add a joint to the scene. From Figure 8, we found that the average total time to create a scene was 25 minutes and 12 seconds of which only 14 minutes and 40 seconds were active work. We also observed that the expert user accomplished the task faster than the rest, and twice as fast as the slowest user. This indicates that with practice, our GUI allows users to become faster at generating scenes. We conclude that doing the real-to-sim transfer of the scenes using the proposed GUI seems to be an intuitive process that is neither time nor labor-intensive when compared to collecting many demonstrations in the real world. We provide more details about the study in Appendix XII.
Figure 8: 3D reconstruction GUI’s user study breakdown times. On average it takes 14 minutes and 40 seconds of active time or 25 minutes and 12 seconds of total time to create a scene through our proposed pipeline.
VI Limitations and Conclusion
Limitations
While our use of 3D point clouds instead of RGB enables easier sim-to-real transfer, we require accurate depth sensors that can struggle to detect thin, transparent, and reflective objects. Future work may investigate applying RialTo to train policies that operate on RGB images or RGBD, as our framework makes no fundamental assumptions that prevent using different sensor modalities. We are also limited to training policies for tasks that can be easily simulated and for real-world objects that can be turned into digital assets. Currently, this is primarily limited to articulated rigid bodies, but advancements in simulating and representing deformables should allow our approach to be applied to more challenging objects. Even though we show RialTo works on fast controllers, these are still relatively slow to minimize the sim-to-real gap in dynamics, thereafter there is potential to investigate tasks for which faster controllers are needed. In this work, we consider relatively quasistatic problems, where exact identification of physics parameters is not necessary for the constructed simulation. This will become important as more complex environments are encountered. Finally, as we explain in Section XIII, RialTo currently takes around 2 days of wall-clock time end-to-end to train a policy for each task, this time bottleneck makes continual learning infeasible and understanding how to obtain policies faster with minimal human supervision would be valuable. We expect with more efficient techniques for learning with point clouds and better parallelization, this procedure can be sped up significantly.
Conclusion
This work presents RialTo, a system for acquiring policies that are robust to environmental variations and disturbances on real-world deployment. Our system achieves robustness through the complementary strengths of real-world imitation learning and large-scale RL on digital twin simulations constructed on the fly. Our results show that by importing 3-D reconstructions of real scenes into simulation and collecting a small amount of demonstration data, non-expert users can program manipulation controllers that succeed under challenging conditions with minimal human effort, showing enhanced levels of robustness and generalization.
Acknowledgments
The authors would like to thank the Improbable AI Lab and the WEIRD Lab members for their valuable feedback and support in developing this project. In particular, we would like to acknowledge Antonia Bronars and Jacob Berg for helpful suggestions on improving the clarity of the manuscript, and Marius Memmel for providing valuable insights on learning from point clouds in the early stages of the project. This work was partly supported by the Sony Research Award, the US Government, and Hyundai Motor Company.
Author Contributions
Marcel Torne conceived the overall project goals, investigated how to obtain real-to-sim transfer of scenes and policies, and robustly do sim-to-real transfer of policies, wrote all the code for the policy learning pipeline and RialTo’s GUI for real-to-sim transfer of scenes, ran simulation and real-world experiments, wrote the paper, and was the primary author of the paper.
Anthony Simeonov helped with setting up the robot hardware, made technical suggestions on learning policies from point clouds, helped with the task of placing the plate on the rack, and actively helped with writing the paper.
Zechu Li assisted in the early stage of conceiving the project and helped develop RialTo’s GUI for the real-to-sim transfer of the scenes.
April Chan led the user study experiments to analyze RialTo’s GUI.
Tao Chen provided valuable insights and recommendations on sim-to-real transfer.
Abhishek Gupta was involved in conceiving the goals of the project, assisted with finding the scope of the paper, suggested baselines and ablations, played an active role in writing the paper and co-advised the project.
Pulkit Agrawal suggested the idea of doing real-to-sim transfer of scenes, was involved in conceiving the goals of the project, suggested baselines and ablations, helped edit the paper, and co-advised the project.